40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 17039     Entry time: Wed Jul 27 14:39:04 2022
Author: Deeksha 
Type: Update 
Category: Electronics 
Subject: New and improved PZT TF data from the DFD 

Paco and I messed around with the attenuation of the scope and bandwidth of the IF. We also replaced the BNC T's in the circuits with RF splitters. We saw some decent improvements to the data. The data is attached and a diagram of the experiment. [We analytically calculated the impedances to avoid any mismatch taking place]. Working on fitting the data.

We also moved around the wires so that the AG4395 is closer to the PZT.

 

Attachment 1: tf_data.png  104 kB  Uploaded Wed Jul 27 15:49:07 2022  | Hide | Hide all
tf_data.png
Attachment 2: latest_data.zip  44 kB  Uploaded Wed Jul 27 18:13:52 2022
Attachment 3: dfd_-_new.drawio.png  57 kB  Uploaded Fri Jul 29 16:16:53 2022  | Hide | Hide all
dfd_-_new.drawio.png
ELOG V3.1.3-